# Fairness-Accuracy Trade-Offs: A Causal Perspective

### v1

@@ Appendix C Connection to Causal Decompositions
$\text{TVD}(S\to S^{\prime})=\underbrace{\mathbbm{E}[Y^{S^{\prime}}\mid x_{1}]-\mathbbm{E}[Y^{S^{\prime}}\mid x_{0}]}_{\text{TV after removing }S^{\prime}\setminus S}-\underbrace{\mathbbm{E}[Y^{S}\mid x_{1}]-\mathbbm{E}[Y^{S}\mid x_{0}]}_{\text{TV before removing }S^{\prime}\setminus S}.$ (78)
In words, the quantity $\text{TVD}(S\to S^{\prime})$ measures how much the TV measure changes once the effect in $S^{\prime}\setminus S$ is constrained to 00. Put differently, the $\text{TVD}(S\to S^{\prime})$ is a quantification of the amount of disparity that is transmitted along the path $S^{\prime}\setminus S$. Other notions for quantifying the size of the disparity transmitted along a specific causal pathway have also been explored in the literature on causal decompositions [Zhang and Bareinboim, 2018b, Plečko and Bareinboim, 2024]. An important result from [Plečko and Bareinboim, 2024] we draw a connection to is the decomposition of the TV measure into the NDE, NIE, and NSE contributions:
Proposition 4 (TV Decomposition [Plečko and Bareinboim, 2024]).
@@
$\text{TV}_{x_{0},x_{1}}(y)=\text{NDE}_{x_{0},x_{1}}(y)-\text{NIE}_{x_{1},x_{0}}(y)+\text{NSE}_{x_{1}}(y)-\text{NSE}_{x_{0}}(y).$ (79)
@@
(a) TVD and TVR comparison on Census 2018 data.
(b) TVD and TVR comparison on COMPAS data.
@@ Appendix D CFUR Quantity Representation
The ATVD and APSEL quantities are defined as the average values of TVD, PSEL quantities, defined through Eqs. 24-25. In this appendix, we discuss the connection of the CFUR value with the local fairness-utility ratios, defined via:
@@
Appendix E UCI Credit Example
@@
Example 4 (UCI Credit [Yeh, 2016]).
A commercial bank is developing an automated payment default scoring system based on the UCI Credit Dataset, collected in 2018. This dataset includes demographic variables $Z$ ($Z_{1}$ for age), the protected attribute gender $X$ ($x_{0}$ female, $x_{1}$ male), marital status $M$, education level $L$, employment status $E$, and financial history $F$. The institution aims to predict the binary outcome $Y$ that indicates whether the individual defaulted on a credit payment ($Y=1$ for defaulting, $Y=0$ otherwise). The grouping $\{X={X},Z=\{Z_{1}\},W=\{L,M,E,F\},Y=Y\}$ induces the standard fairness model [Plečko and Bareinboim, 2024] for this analysis.

### v2

@@ Appendix C Connection to Causal Decompositions
(a) TVD and TVR comparison on Census 2018 data.
(b) TVD and TVR comparison on COMPAS data.
@@
$\text{TVD}(S\to S^{\prime})=$ $\underbrace{\mathds{E}[Y^{S^{\prime}}\mid x_{1}]-\mathds{E}[Y^{S^{\prime}}\mid x_{0}]}_{\text{TV after removing }S^{\prime}\setminus S}$ (102)
$-\underbrace{\mathds{E}[Y^{S}\mid x_{1}]-\mathds{E}[Y^{S}\mid x_{0}]}_{\text{TV before removing }S^{\prime}\setminus S}.$ (103)
In words, the quantity $\text{TVD}(S\to S^{\prime})$ measures how much the TV measure changes once the effect in $S^{\prime}\setminus S$ is constrained to 00. Put differently, the $\text{TVD}(S\to S^{\prime})$ is a quantification of the amount of disparity that is transmitted along the path $S^{\prime}\setminus S$. Other notions for quantifying the size of the disparity transmitted along a specific causal pathway have also been explored in the literature on causal decompositions (Zhang and Bareinboim 2018b; Plečko and Bareinboim 2024). An important result from (Plečko and Bareinboim 2024) we draw a connection to is the decomposition of the TV measure into the NDE, NIE, and NSE contributions:
Proposition 4 (TV Decomposition (Plečko and Bareinboim 2024)).
@@
$\text{TV}_{x_{0},x_{1}}(y)$ $=\text{NDE}_{x_{0},x_{1}}(y)-\text{NIE}_{x_{1},x_{0}}(y)$ (104)
$\quad+\text{NSE}_{x_{1}}(y)-\text{NSE}_{x_{0}}(y).$ (105)
@@ Appendix D CFUR Quantity Representation
The ATVD and APSEL quantities are defined as the average values of TVD, PSEL quantities, defined through Eqs. 40-41. In this appendix, we discuss the connection of the CFUR value with the local fairness-utility ratios, defined via:
@@
Appendix E Additional Experiments
In this appendix, we analyze the COMPAS dataset (Larson et al. 2016) (Ex. 3), and the UCI Credit dataset (Yeh 2016) (Ex. 4).
(a) (A)PSEL and (A)TVD values.
(b) CFUR values.
(c) Fairness/Utility of $\widehat{Y}^{S}$ predictors.
(d) $\mathcal{G}_{\text{PSEL}}$ with PSEL (blue) and TVD values (red).
Figure 7: Application of Alg. 1 on the COMPAS dataset.
Example 3 (Recidivism Prediction on COMPAS (Larson et al. 2016)).
The team from ProPublica wishes to understand the tension between fairness and accuracy, since the latter is an important concern of the district court. For the analysis, they use the complement of the area under the receiver operator characteristic curve ($1-$AUROC). They apply Alg. 1, and find that the APSEL value is the largest for the indirect effect (around 12% of AUROC), while it is smaller for the spurious effect (6%), and negligible for the direct effect, implying that the direct effect does not play a significant role in the prediction (see Fig. 7(a)). Regarding TVD values, the team finds that the removal of the spurious effect results in a 3% decrease in the TV measure, while the removal of the indirect effect decreases the disparity by 5%. Based on this, they find that the spurious effect is the best for reducing the disparity between groups (see Fig. 7(b) for CFUR values). Finally, the team also visualizes the TV and excess loss dependence (Fig. 7(c)), and the graph $\mathcal{G}_{\text{PSEL}}$ with PSEL and TVD values associated with each transition and effect removal (Fig. 7(a)). In the upcoming court hearing, the team proposes the usage of the causally-constrained predictor $\widehat{Y}^{\{D,S\}}$ that removes direct and spurious effects, and they use this analysis to demonstrate the impact of this choice on accuracy.
@@ Appendix E Additional Experiments
Example 4 (UCI Credit (Yeh 2016)).
A commercial bank is developing an automated payment default scoring system based on the UCI Credit Dataset, collected in 2018. This dataset includes demographic variables $Z$ ($Z_{1}$ for age), the protected attribute gender $X$ ($x_{0}$ female, $x_{1}$ male), marital status $M$, education level $L$, employment status $E$, and financial history $F$. The institution aims to predict the binary outcome $Y$ that indicates whether the individual defaulted on a credit payment ($Y=1$ for defaulting, $Y=0$ otherwise). The grouping $\{X={X},Z=\{Z_{1}\},W=\{L,M,E,F\},Y=Y\}$ induces the standard fairness model (Plečko and Bareinboim 2024) for this analysis.
@@
Appendix F Path-Specific Fairness-Accuracy Trade-Offs
In this appendix, we translate the results from the main text to a more general setting with path-specific causal effects. In particular, we consider a setting in which $m$ distinct paths are considered, labeled $S_{1},\dots,S_{m}$. A particular choice we may consider is to set $S_{1}$ to measure the direct effect, $S_{2}$ to $S_{k+1}$ to measure the contributions of each of the confounders $Z_{1},\dots,Z_{k}$ to the overall spurious effect, and $S_{k+2}$ to $S_{k+\ell+1}$ to measure the contributions of each of the mediators $W_{1},\dots,W_{\ell}$ to the overall indirect effect. For explicit expressions for $S_{1},\dots,S_{m}$ we refer the reader to a detailed discussion in (Plecko and Bareinboim 2024a, Sec. 6), where some options for instantiating the effects $S_{1},\dots,S_{m}$ are introduced and explained.
We now go over the key results of the main paper that need to be adapted, and explain how they change when we consider path-specific effects. A causally-fair predictor for a subset $S\subseteq\{S_{1},\dots,S_{m}\}$ of the above measures is defined as the following solution:
Definition 6 (Path-Specific Causally Fair Predictor (Plecko and Bareinboim 2024b)).
$\operatorname*{arg\,min}_{f}$ $\mathds{E}\;L(Y,f(X,Z,W))$ (122)
s.t. $S_{i}(f)=S_{i}(y)\mathbb{1}(S_{i}\notin S)\quad\forall i.$ (123)
The above definition is a generalization of Def. 2. Building on this definition, we can state a more general version of Thm. 1:
Theorem 6 (Total Excess Loss Path-Specific Decomposition).
The total excess loss $\text{PSEL}(\emptyset\to\{S_{1},\dots,S_{m}\})$ can be decomposed into a sum of path-specific excess losses as follows:
$\text{PSEL}(\emptyset\to\{S_{1},\dots,S_{m}\})$ $=\sum_{i=0}^{m-1}\text{PSEL}(S_{[i]}\to S_{[i+1]}),$ (124)
where $S_{[i]}$ is the set $S_{1},\dots,S_{i}$, and $S_{[0]}=\emptyset$.
For Algs. 1 and  2, we also need some modifications. Crucially, in Alg. 1, we compute the APSEL$(S_{i})$ values as
$\frac{1}{m!}\sum_{\text{permutations }\sigma}\text{PSEL}(\tilde{S}_{[\sigma^{-1}(i)]}\to\tilde{S}_{[\sigma^{-1}(i)+1]}),$ (125)
and the $\text{ATVD}(S_{i})$ value as
$\frac{1}{m!}\sum_{\text{permutations }\sigma}\text{TVD}(\tilde{S}_{[\sigma^{-1}(i)]}\to\tilde{S}_{[\sigma^{-1}(i)+1]}),$ (126)
where $\sigma^{-1}(i)$ represents the position of index $i$ in the permutation, and $\tilde{S}$ is the permuted ordering of $S_{1},\dots,S_{m}$ according to $\sigma$. The above summations range over all $m!$ permutations $\sigma$.
Now, for adapting Alg. 2, we first provide a definition of a path-specific Lagrangian causally-fair predictor:
Definition 7 (Path-Specific Causal Lagrange Predictor $\widehat{Y}^{S}$).
The causally $S$-fair $\lambda$-optimal predictor $\widehat{Y}^{S}(\lambda)$ with respect to pathways in $S\subseteq\{S_{1},\dots,S_{m}\}$ and the loss function $L$ is the solution to the following optimization problem:
$\operatorname*{arg\,min}_{f}$ $\;\mathds{E}\;L(Y,f(X,Z,W))+$ (127)
$\lambda\sum_{i=1}^{m}\Big{(}S_{i}(f)-S_{i}(y)\mathbb{1}(S_{i}\notin S)\Big{)}^{2}$ (128)
This definition builds on the notion introduced in Def. 2, but considers a possibly broader range of causal pathways. We also present Alg. 3, which is an appropriate generalization of Alg. 2 for the more general path-specific case.
Input: training data $\mathcal{D}_{t}$, evaluation data $\mathcal{D}_{e}$, set $S$, precision $\epsilon$
1 Set $\lambda_{\text{low}}=0$, $\lambda_{\text{high}}=$ large
2 while $|\lambda_{\text{high}}-\lambda_{\text{low}}|>\epsilon$ do
4       fit a neural network to solves the optimization problem in Eqs. 127-128 with $\lambda=\lambda_{\text{mid}}$ on $\mathcal{D}_{t}$ to obtain the predictor $\widehat{Y}^{S}(\lambda_{\text{mid}})$
5       compute the causal measures of fairness $S_{1},\dots,S_{m}$ for $\widehat{Y}^{S}(\lambda_{\text{mid}})$ on evaluation data $\mathcal{D}_{e}$
6       test the hypothesis
7       for $i\in{1,\dots,m}$ do
8
$H_{0}^{S_{i}}:S_{i}(\widehat{y}^{S}(\lambda_{\text{mid}}))$ $=S_{i}(y)\cdot\mathbb{1}(S_{i}\notin S)$ (129)
9      if any $H_{0}^{S_{i}}$ rejected then $\lambda_{\text{low}}=\lambda_{\text{mid}}$ else $\lambda_{\text{high}}=\lambda_{\text{mid}}$
return predictor $\widehat{Y}^{S}(\lambda_{\text{mid}})$
Algorithm 3 Causally-Fair Constrained Learning (CFCL)
In Alg. 3, $m$ causal pathways are considered, and we perform a larger set of hypothesis tests when determining the appropriate value of the tuning parameter $\lambda$. In particular, in each step of the binary search procedure in Alg. 3, we test all of the hypotheses:
$H_{0}^{S_{i}}:S_{i}(\widehat{y}^{S}(\lambda_{\text{mid}}))$ $=S_{i}(y)\cdot\mathbb{1}(S_{i}\notin S)$ (130)
on the evaluation data $\mathcal{D}_{e}$. In this way, we verify if each of the effects $S_{i}$ takes an appropriate value on a held-out dataset, which ensures that the correct path-specific constraints hold on the held-out data.
Further, we note that the statement of Prop. 3 remains the same, and the proof of the more general statement considering $m$ causal pathways is given in Appendix B (see proof of Prop. 3). With this, the full set of results in the main paper are described for the setting with path-specific effects $S_{1},\dots,S_{m}$.